Let a consumer’s utility function be π‘ˆ = π‘ž1

6π‘ž2

4 + 1.5πΌπ‘›π‘ž1 + πΌπ‘›π‘ž2

and his budget constraint

3π‘ž1 + 4π‘ž2 = 100.

a. Find his optimum commodity bundle of π‘ž1

and π‘ž2

.Β 

b. Verify whether second order is fulfilled or not.Β 

c. Determine the marginal utility of money

d. Estimate the new optimal utility if the consumer’s income rises by ₦1


1
Expert's answer
2021-03-02T07:46:32-0500
"\frac{\delta U}{\delta q_1}=6q_1^5q_2^4+\frac{1.5}{q_1}"

"\frac {\delta U}{\delta q_2}=4q_1^6q_2^3+\frac{1}{q_2}"

"\frac{6q_1^5q_2^4+\frac{1.5}{q_1}}{3}=\frac {4q_1^6q_2^3+\frac{1}{q_2}}{4}"


"q_1=20"

"q_2=10"

"3\times10+4\times20=100"

"\lambda=9.6\times10^{10}"

If I=101

Since the growth in consumer income is relatively small compared to previous income, the Lagrange multiplier will not change significantly.